# Computing finite–temperature elastic constants with noise cancellation

### v1

@@ Abstract
Elastic constants are central material properties, frequently reported in experimental and theoretical studies. While their computation is straightforward in the absence of thermal fluctuations, finite–temperature methods often suffer from poor signal–to–noise ratios or the presence of strong anharmonic effects. Here, we show how to compute elastic constants in thermal ordered and disordered systems by generalizing a noise–cancellation method originally developed for piezoelectric coupling coefficients. A slight strain is applied to an equilibrated solid. Simulations of both the strained and unstrained (or oppositely strained) reference systems are performed using identical thermostatting schemes. As demonstrated theoretically and with generic one–dimensional models, this allows stress differences to be evaluated and elastic constants to be determined with much reduced thermal noise. We then apply this approach across a diverse set of systems, spanning crystalline argon, ordered silicon as well as amorphous silicon, poly(methyl methacrylate), and cellulose derivatives.
@@ I Introduction
Elastic constants are routinely computed in molecular simulations using two mainstream methods: response–to–deformation (RTD) [1] and fluctuation–based (FB) [2, 3] computations. Their central ideas are as follows: When deforming a solid from a stress–free state, the strain tensor elements, $\varepsilon_{i}$ in Voigt notation, are known, and the resulting stress–tensor elements, $\sigma_{i}$, can be computed. The elastic tensor elements, $C_{ij}$, then follow from $\sigma_{i}=C_{ij}\varepsilon_{j}$. Alternatively, the simulation box can be fixed, and the covariance of the stress–tensor elements can be related directly to the elastic tensor: $C_{ij}=V\langle\delta\sigma_{i}\delta\sigma_{j}\rangle/(k_{\rm B}T)$, where $V$ is the volume of the simulation box and $k_{\rm B}T$ is the thermal energy, with $k_{\rm B}$ being the Boltzmann constant. Analyzing box–shape fluctuations in constant–stress simulations is also a viable approach [2] provided that meaningful barostats are employed, which can be difficult to find in standard software. Worth mentioning is also the determination of elastic constants through the computation of the sound velocity [4, 5], which can be regarded as a FB method. This approach, however, relies on analyzing dynamical rather than static covariances.
@@
Each of the mainstream methods includes further sub–methods [6]. For example, at zero temperature, it suffices to compute the energy as a function of strain, though non–affine displacements may have to be accounted for. Alternatively, one can compute second-order derivatives of the energy with respect to atomic coordinates, where self–affine displacements can again cause complications [6, 7]. Additional modifications apply in the presence of thermal fluctuations [8, 9] or when the reference is stressed [7]. Accounting for quantum fluctuations, not considered here, requires yet additional terms [10, 11].
@@
RTD and FB computations of elastic tensor elements each have their own advantages and disadvantages. In particular, FB computations require long convergence times; using larger system sizes even increases the stochastic error, given fixed compute time, while the RTD method would leave it constant. However, RTD computations conducted at what may deceptively be considered a small strain, say, 1%, can produce errors in the elastic constants exceeding 10% due to nonlinear effects, whereas simulations at smaller strains are dominated by stochastic noise.
@@
Here, we extend a noise-cancellation technique originally developed for the piezoelectric coefficient $\gamma$ of $\alpha$–quartz, which remained effective even near the $\alpha$$\beta$ quartz transition temperature [12]. The method contrasted the strain in two constant–stress simulations, both initialized from the same equilibrated configuration and both using the same random number sequence in the thermostat. In one sample, an electric field was applied, which slowly increased over time, while the other sample remained field–free. The strain difference between the samples was essentially free of thermal fluctuations– at least during the initial stages of the simulation– and allowed $\gamma$ to be deduced reliably within a few hundred time steps. This approach can be classified as a correlated sampling technique, which is more commonly employed in Monte Carlo than in molecular dynamics simulations, and was potentially first formulated in the context of nuclear physics [13].
@@ II.1 Preliminary considerations
The preliminary considerations discussed above highlight why computing the elastic constants of (amorphous) polymers presents particular challenges. Specifically, the steepness of the Lennard–Jones (LJ)–like interactions necessitates the use of extremely small strain increments in finite–difference (FD)–based schemes for reliable response–to–deformation (RTD) calculations. At the same time, the relatively low energy scales– characterized by small interaction prefactors (e.g., $U_{0}$ typically being below the thermal energy at room temperature)– make the system highly susceptible to thermal fluctuations. These combined effects contribute to significant noise in stress responses and complicate the accurate determination of the elastic tensor components.
@@ II.2 Noise cancellation in ideally harmonic systems
where $k_{s}>k_{l}$ and $a_{s}^{0}<a_{l}^{0}$ are the stiffnesses and equilibrium bond lengths of the short ($s$) and long ($l$) bonds, respectively. The box length per unit cell is denoted by $h$, in analogy to the $h$-matrix of constant–stress simulations [15], and $x_{21}=x_{2}-x_{1}$. Here, $x_{i}$ are reduced coordinates, with real positions $X_{i}=hx_{i}$. They do not appear in the final expressions but are required for intermediate steps. Reduced coordinates are also the natural variables in the constant–stress ensemble, ensuring a direct link between the analytical treatment and simulations. Finally, $\sigma$ is the external (tensile) stress. For readability, the ideal gas contribution is omitted here but will be included where necessary.
@@ II.3 Noise cancellation in quasi–harmonic systems
Eq. (16) indicates that the stress estimators, which are proportional to $\Delta V_{2}/\varepsilon$, begin to deviate at a rate independent of $\varepsilon$. Initially, the deviation between the potential energy in the strained and reference samples grows as $\Delta t^{2}$, as long as $\Delta t$ remains small compared to the relevant phonon periods. Once $\Delta t$ exceeds $1/\omega_{e}(0)$ but is still smaller than $1/\varepsilon\omega_{e}^{\prime}$, the deviation grows linearly in time. However, a linear regime should not be expected in extended systems due to the broad distribution of eigenfrequencies. Because the (instantaneous) stress, and hence the elastic–constant estimator, scale with $\Delta V_{2}/\varepsilon$, the time dependence of the spurious contribution is, to leading order, independent of $\varepsilon$. This contribution is called undesired because it masks the genuine relaxation signal of self–affine modes.
@@
While Figure 2 already confirmed the expectation that larger damping implies smaller standard deviation of the instantaneous measurements, Figure 3 corroborates this result using a double–logarithmic representation of the standard deviation as a function of time. This analysis also shows that the moment in time, when the standard deviation starts to increase does not depend on the strain $\varepsilon$, at least as long as it is sufficiently small. Moreover, it grows to a larger final value as the strain decreases. Yet, it remains far below the stochastic errors that one would obtain if the two follow–up simulations were run using a different set of random numbers. Perhaps, most importantly, when using sufficiently large damping, the long–time standard deviation is independent in leading order on $\varepsilon$.
@@ III.2 Silicon
The crystalline sample is initialized in the diamond cubic structure, using a lattice constant of $a=5.431$ Å [7]. Starting from this configuration, the amorphous sample is generated by heating the system to $T=4,000$ K to induce melting, followed by an immediate quench to $T=300$ K. Both the heating and cooling phases are carried out over 250 ps using a time step of $\Delta t=0.5$ fs. During these stages, a constant pressure of $p=1$ bar is maintained using the Parrinello–Rahman barostat. The elastic tensor components $C_{ij}$ are computed at $T=300$ K, with temperature controlled via a Langevin thermostat with Grønbech–Jensen algorithm [21], using a time constant $\tau_{\rm T}=0.01$ ps. These simulations are typically run for 10 ps with a time step of $\Delta t=10$ fs. In some cases, the convergence of $C_{ij}$ is also tested with different $\tau_{\rm T}$; such details are reported where relevant.
@@ III.5 Computation of elastic tensor elements
$\varepsilon_{2}$ $=$ $\frac{1}{1+\varepsilon}$ (19b)
@@ IV.1 Effect of thermostat
The calculation in Section II.2 was based on the assumption that $C(t)$ in Eq. (10) has decayed. Thermostatting should therefore aim to achieve this decay as efficiently as possible. Therefore, we begin by examining the impact of thermostat choice on simulations of crystalline argon and silicon, as shown in Figure 5 and Figure 6. Interestingly, the widely used canonical sampling through velocity rescaling (CSVR) thermostat [22]– a global kinetic energy control method– produces dynamics that closely resemble those of the microcanonical (NVE) ensemble. A notable difference, however, emerges in the initial relaxation behavior of the shear modulus estimator between thermostatted dynamics (solid black line) and pseudo–microcanonical schemes (dotted blue and dashed red lines). This difference arises because the microcanonical ensemble does not enable energy exchange between vibrational modes, which can limit convergence of the shear modulus, $G$, over time– particularly in nearly harmonic systems.
Despite above discrepancies– between the properly thermostatted and microcanoncal simulations– in transient behavior, the time–averaged values of $G$ differ by less than 1.5% across the different approaches. However, the associated statistical uncertainty– estimated via block–averaged standard deviations– exceeds 10% for both the NVE and CSVR ensembles, while it drops to approximately 0.5% when using the Langevin thermostat [21]. Although similar trends are observed for both crystalline argon at $T=10$ K and silicon at $T=300$ K (both well witin their respective harmonic limits, i.e., significantly below the Debye temperature), the case of silicon is more nuanced due to non–affine displacements of the basis atoms in the diamond structure under anisotropic deformation.
We note in passing why the microcanonical setup exhibits large fluctuations. Ideally, in the absence of external noise, it should yield a well–converged value. However, as shown in Eq. (10), the Green’s function term vanishes, leaving only the first term, which does not decay over time. Importantly, this term inherits fluctuations from the reference system that also do not decay with time, as reflected in the behavior of $C(t)$.
@@ IV.2.1 Interpretation
Figure 8: A summary of the components of the elastic modulus tensor for various systems, as calculated using the proposed method. Notably, the results for crystalline argon and silicon are averaged over 1 and 16 independent runs, respectively, while the amorphous silicon data represent an average over 49 runs. For all polymer systems, the reported values are averaged over 100 independent simulations. The polymer data shown herein have been coarse–grained for clarity.
@@
IV.2.2 Application
Figure 7 presents results for two different samples, each simulated using a range of Langevin damping times $\tau_{\rm T}$. At small damping times– where stiff vibrational modes are critically damped– they equilibrate rapidly. However, this also leads to a slower overall relaxation of $C_{11}$. In contrast, larger damping times result in slower equilibration of high–frequency modes, but the average value around which they fluctuate converges more quickly to the long–time relaxed value of $C_{11}$. A trend that is also evident for a simplified model in Figure 2.
@@ IV.2.2 Application
For the production runs of all systems studied in this work, we selected damping times, $\tau_{\rm T}$, that ensure rapid convergence of the elastic constants, $C_{ij}$, to their reference values. Added advantage of the single damping time scheme is that it is easy to implement even in the complex polymeric systems. Figure 8 summarizes the time evolution and convergence behavior of $C_{ij}$ for each system. Notably, after the initial rapid relaxation, the amorphous samples exhibit a secondary, slower decay in the elastic constants $C_{ij}$ over time. This effect becomes increasingly pronounced with greater structural complexity– for instance, the decay is noticeably slower in amorphous cellulose compared to amorphous silicon (see Figures 8e, k, and n). A similar two–stage relaxation trend is observed in the case of the generic diatomic harmonic chain with thermal fluctuations (see Figure 1). At first glance, and in the absence of the theoretical insights provided by Figure 1, this transient behavior might be misinterpreted as an indication of flow or irreversible deformation within the sample. To rule out this possibility, we conducted a series of additional simulations involving energy minimization on pre–strained cellulose samples. In these tests, the atomic coordinates remained unchanged between the strained configurations and the corresponding energy–minimized states (data not shown), thereby confirming the absence of any flow–like behavior. Taken together, these observations suggest that the slow decay in $C_{ij}$ is not due to structural rearrangements but rather arises from intrinsic vibrational eigenfrequencies of the system. Specifically, the broad distribution of eigenfrequencies characteristic of polymeric and amorphous materials leads to a slow convergence of the elastic response. In particular, the overpopulation of low–frequency vibrational modes– commonly referred to as the Boson peak [33]– is a key factor underlying the long-time decay of $C_{ij}$.
Table 1: A table listing the average values of the elastic modulus components. The values are computed by block–averaging over ten blocks taken from the second half of the data shown in Figure 8. The error bars represent the standard deviations across these ten blocks. For comparison, available literature data for different systems are also included.
@@
Argon This work $0.868\pm 0.002$ $4.069\pm 0.007$ $2.322\pm 0.003$
@@
Experiment [35] 4.0 2.0
Silicon This work $42.8\pm 0.2$ $154.8\pm 0.8$ $69.8\pm 0.9$
Previous sim. [23] 70 150 80
@@
Silicon This work $23.0\pm 1.6$ $113.0\pm 1.3$ $66.6\pm 1.1$
@@
PMMA This work $1.9\pm 0.6$ $6.4\pm 0.5$ $3.1\pm 0.6$
@@
Cellulose This work $2.7\pm 0.2$ $14.8\pm 0.2$ $8.6\pm 0.3$
Cellulose Acetate This work $2.1\pm 0.1$ $10.9\pm 0.1$ $6.6\pm 0.1$
A summary of the averaged $C_{ij}$ values, along with available literature data for comparison, is provided in Supplementary Table 1. The results exhibit good agreement with previously published values, thereby supporting the consistency and reliability of our computational approach. Notably, the values reported for cellulose and cellulose acetate are predictive in nature, as– to the best of our knowledge– no corresponding $C_{ij}$ data are currently available in the simulation or experimental literature. These predictions provide a valuable reference point for future studies. Furthermore, the computed $C_{ij}$ values for cellulose derivatives across a range of temperatures have already been successfully employed to estimate thermal transport coefficients $\kappa$ in related work [29], which was the main motivation behind the development of the noise cancellation approach proposed herein. Lastly, we emphasize that ideal gas corrections to the computed $C_{ij}$ values are negligible in all cases, remaining below 0.1%, and therefore do not influence the reported results.

### v2

@@ Abstract
Elastic constants are central material properties, frequently reported in experimental and theoretical studies. While their computation is straightforward in the absence of thermal fluctuations, finite–temperature methods often suffer from poor signal–to–noise ratios, the presence of strong anharmonic effects, or require second-order derivatives with respect to spatial coordinates to be taken. Here, we show how to compute elastic constants in thermal ordered and disordered systems by generalizing a noise–cancellation method originally developed for piezoelectric coupling coefficients. A slight strain is applied to an equilibrated solid. Simulations of both the strained and unstrained (or oppositely strained) reference systems are performed using identical thermostatting schemes. As demonstrated theoretically and with generic one–dimensional models, this allows stress differences to be evaluated and elastic constants to be determined with favorable thermal noise. We then apply this approach across a diverse set of systems, spanning crystalline argon, ordered silicon as well as amorphous silicon, poly(methyl methacrylate), and cellulose derivatives.
@@ I Introduction
Elastic constants are routinely computed in molecular simulations using two mainstream methods: response–to–deformation (RTD) [1] and fluctuation–based (FB) [2, 3] computations. Their central ideas are as follows: When deforming a solid from a stress–free state, the strain tensor elements, $\varepsilon_{i}$ in Voigt notation, are known, and the resulting stress–tensor elements, $\sigma_{i}$, can be computed. The elastic tensor elements, $C_{ij}$, then follow from $\sigma_{i}=C_{ij}\varepsilon_{j}$. Alternatively, the simulation box can be fixed, and an estimator be evaluated, which consists of a so-called Born term and a stress-fluctuation correction [4, 5]. The Born term hinges on the Hessian matrix and is thus frequently not available, in particular for many-body or machine-learned potentials. Analyzing box–shape fluctuations in constant–stress simulations is also a viable approach [2] provided that meaningful barostats are employed, which can be difficult to find in standard software. Worth mentioning is also the determination of elastic constants through the computation of the sound velocity [6, 7], which can be regarded as a FB method. This approach, however, relies on analyzing dynamical rather than static covariances.
@@
Each of the mainstream methods includes further sub–methods [8, 9]. For example, at zero temperature it suffices to compute the energy as a function of strain, although internal modes must be allowed to relax in a non–affine fashion if applicable [1, 9]. Alternatively, one can compute second–order derivatives of the energy with respect to atomic coordinates, where non–affine displacements can again cause complications [1, 9], in particular for a stressed reference [10]. Additional modifications apply in the presence of thermal fluctuations [4, 11, 5] or, again, when the reference is stressed [10]. Accounting for quantum fluctuations, not considered here, requires additional terms [12, 13].
@@
RTD and FB computations of elastic tensor elements each have their own advantages and disadvantages. In particular, FB computations require long convergence times; using larger system sizes even increases the stochastic error, given fixed compute time, while the RTD method would leave it constant. However, RTD computations using simple forward-difference schemes and conducted at what may deceptively be considered a small strain, say, 1%, can produce errors in the elastic constants on the order of 10% due to nonlinear effects, whereas simulations at smaller strains are dominated by stochastic noise.
An alternative to the above–mentioned algorithms is based on the diagonalization of the Hessian matrix while accounting for non–affine deformations in the system. Although such approaches are computationally expensive, particularly for large system sizes, they yield exact results in the absence of thermal fluctuations and allow local elastic constants to be defined [14, 15]. When the temperature is finite, anharmonic effects set in, which can be captured by considering the Hessian of the instantaneous rather than the relaxed normal modes [16]. However, the occurrence of negative eigenvalues in the Hessian makes a rigorous error estimate difficult. Even the computation of viscoelastic properties of ultra-strong disordered systems, which lack barrier crossings, is complex when done with care [17]. In this context, the present work focuses on a lean computation of only the elastic tensor, prioritizing minimal implementation, low noise, and minimal systematic errors.
@@
Here, we extend a noise–cancellation technique originally developed for the piezoelectric coefficient $\gamma$ of $\alpha$–quartz, which remained effective even near the $\alpha$$\beta$ quartz transition temperature [18]. The method contrasted the strain in two constant–stress simulations, both initialized from the same equilibrated configuration and both using the same random number sequence in the thermostat. In one sample, an electric field was applied, which slowly increased over time, while the other sample remained field–free. The strain difference between the samples was essentially free of thermal fluctuations– at least during the initial stages of the simulation– and allowed $\gamma$ to be deduced reliably within a few hundred time steps. This approach can be classified as a correlated sampling technique, which is more commonly employed in Monte Carlo than in molecular dynamics simulations, and was potentially first formulated in the context of nuclear physics [19].
@@ II.1 Preliminary considerations
The preliminary considerations discussed above highlight why computing the elastic constants of (amorphous) polymers presents particular challenges. Specifically, the steepness of the Lennard–Jones (LJ)–like interactions necessitates the use of very small strain increments in finite–difference (FD)–based schemes for reliable response–to–deformation (RTD) calculations. At the same time, the relatively low energy scales– characterized by small interaction prefactors (e.g., $U_{0}$ typically being below the thermal energy at room temperature)– make the system highly susceptible to thermal fluctuations. These combined effects contribute to significant noise in stress responses and complicate the accurate determination of the elastic tensor components.
@@ II.2 Noise cancellation in ideally harmonic systems
where $k_{s}>k_{l}$ and $a_{s}^{0}<a_{l}^{0}$ are the stiffnesses and equilibrium bond lengths of the short ($s$) and long ($l$) bonds, respectively. The box length per unit cell is denoted by $h$, in analogy to the $h$-matrix of constant–stress simulations [21], and $x_{21}=x_{2}-x_{1}$. Here, $x_{i}$ are reduced coordinates, with real positions $X_{i}=hx_{i}$. They do not appear in the final expressions but are required for intermediate steps. Reduced coordinates are also the natural variables in the constant–stress ensemble, ensuring a direct link between the analytical treatment and simulations. Finally, $\sigma$ is the external (tensile) stress. For readability, the kinetic-energy contribution is omitted here but will be included where necessary.
@@ II.3 Noise cancellation in quasi–harmonic systems
Eq. (16) indicates that the stress estimators, which are proportional to $\Delta V_{2}/\varepsilon$, begin to deviate at a rate independent of $\varepsilon$. Initially, the deviation between the potential energy in the strained and reference samples grows as $\Delta t^{2}$, as long as $\Delta t$ remains small compared to the relevant phonon periods. Once $\Delta t$ exceeds $1/\omega_{e}(0)$ but is still smaller than $1/\varepsilon\omega_{e}^{\prime}$, the deviation grows linearly in time. However, a linear regime should not be expected in extended systems due to the broad distribution of eigenfrequencies. Because the (instantaneous) stress, and hence the elastic–constant estimator, scale with $\Delta V_{2}/\varepsilon$, the time dependence of the spurious contribution is, to leading order, independent of $\varepsilon$. This contribution is disadvantageous because it masks the genuine relaxation signal of self–affine modes.
@@
While Figure 2 already confirmed the expectation that larger damping implies smaller standard deviation of the instantaneous measurements, Figure 3 corroborates this result using a double–logarithmic representation of the standard deviation as a function of time. This analysis also shows that the moment in time, when the standard deviation starts to increase does not depend on the strain $\varepsilon$, at least as long as it is sufficiently small. Moreover, after transiting through another regime, the std levels off at a larger final value as the strain decreases. Despite being large, the std remains far below the stochastic errors that one would obtain if the two follow–up simulations were run using a different set of random numbers. Perhaps, most importantly, when using sufficiently large damping, the long–time standard deviation is independent in leading order on $\varepsilon$.
@@
Fig. 3 reveals that standard deviations can be large with non-optimized parameter choices. We thus explored the performance compared to the ‘traditional’ method using the Born term and a stress-fluctuation correction [5] for the linear LJ chain. We found a wide spread of results, with standard deviations ranging from being on par to being reduced by a factor of 100 when using the noise cancellation technique. However, while Born terms are readily evaluated for pair potentials, this is not true for many other potentials. Thus, the important message of Fig. 3 is that the noise cancellation technique yields smaller stochastic errors than uncorrelated differences.
@@ III.2 Silicon
The crystalline sample is initialized in the diamond cubic structure, using a lattice constant of $a=5.431$ Å [10]. Starting from this configuration, the amorphous melt is generated by heating the system to T=4, [^0] T=4{,}000 K to induce melting. The resulting melt is then quenched to 300 K at a cooling rate of 1.48 K/ps using $\Delta t=5$ fs and $\gamma_{\rm T}=1.0$ ps. During this stage, a constant pressure of $p=1$ bar is maintained using the Parrinello–Rahman barostat. The elastic tensor components $C_{ij}$ are computed at $T=300$ K, with temperature controlled via a Langevin thermostat with Grønbech–Jensen algorithm [27], using a time constant $\tau_{\rm T}=0.01$ ps. These simulations are typically run for 10 ps with a time step of $\Delta t=1$ fs. In some cases, the convergence of $C_{ij}$ is also tested with different $\tau_{\rm T}$; such details are reported where relevant. We have also tested alternative quenching protocols for preparing the amorphous silicon glass. The corresponding details are provided in Supplementary Section S1 [31].
@@ III.5 Computation of elastic tensor elements
$\varepsilon_{2}$ $=$ $\frac{1}{1+\varepsilon}-1$ (19b)
@@ IV.1 Effect of thermostat
The calculation in Section II.2 was based on the assumption that $C(t)$ in Eq. (10) has decayed. Thermostatting should therefore aim to achieve this decay as efficiently as possible. Therefore, we begin by examining the impact of thermostat choice on simulations of crystalline argon and silicon, as shown in Figure 5 and Figure 6. Interestingly, the widely used canonical sampling through velocity rescaling (CSVR) thermostat [28]– a global kinetic energy control method– produces dynamics that closely resemble those of the microcanonical (NVE) ensemble. While all schemes appear to level off at a similar estimate for the shear modulus at very early times, a notable difference, however, emerges in the later relaxation between thermostatted dynamics (solid black line) and pseudo–microcanonical schemes (dotted blue and dashed red lines). This difference arises because the microcanonical ensemble does not enable energy exchange between vibrational modes, which can limit convergence of the shear modulus, $G$, over time– particularly in nearly harmonic systems. Increasing the system size reduces the stochastic noise. Additional simulations on crystalline silicon, in which the system size was increased from $N=1,000$ to 8,000, suggest that this reduction obeys the usual curse of stochastic sampling, i.e., that system sizes have to be quadrupled to double the signal-to-noise ratio (data not shown).
Despite above discrepancies– between the properly thermostatted and microcanoncal simulations– in transient behavior, the time–averaged values of $G$ differ by less than 1.5% across the different approaches. However, the associated statistical uncertainty– estimated via block–averaged standard deviations– exceeds 10% for both the NVE and CSVR ensembles, while it drops to approximately 0.5% when using the Langevin thermostat [27]. Although similar trends are observed for both crystalline argon at $T=10$ K and silicon at $T=300$ K (both well witin their respective harmonic limits, i.e., significantly below the Debye temperature), the case of silicon is more nuanced due to non–affine displacements of the basis atoms in the diamond structure under anisotropic deformation. Note that the quantum flcutuations only alter elastic constants well below the Debye temperature [12, 39, 40].
We note in passing why the microcanonical setup exhibits large fluctuations. Ideally, in the absence of external noise, it should yield a well–converged value. However, as shown in Eq. (10), the Green’s function term, $g(t)$, vanishes, leaving only the first term, which barely decays over time without explicit damping.
In addition to the system-size dependence of the signal-to-noise ratio, one may be concerned about finite-size effects in the signal itself. For crystalline systems, this is typically minor, as sufficiently large simulation cells effectively provide a “$k$-point sampling” in the sense of the trapezoidal rule and the response is dominated by the $k=0$ contribution. In macroscopically homogeneous but microscopically disordered systems, more pronounced finite-size effects must be expected. In practice, corrections will become small once the linear simulation cell size exceeds at least twice the structural (Ornstein–Zernike) correlation length (typically around $1.5$ nm), and, in the case of polymers, the corresponding end–to–end distance. With linear box sizes of 7.8 nm for PMMA [33], 18.0 nm for cellulose, and 21.6 nm for cellulose acetate [36], our simulations satisfy these criteria comfortably so that finite-size effects are expected to remain small.
@@ IV.2 Effect of thermostat damping time
Figure 8: Notably, the results for crystalline argon and silicon are averaged over 1 and 16 independent runs, respectively, while the amorphous silicon data represent an average over 49 runs. For all polymer systems, the reported values are averaged over 100 independent simulations. The polymer data shown herein have been coarse–grained for clarity.
@@
IV.3 Applications
Figure 7 presents results for two different samples, each simulated using a range of Langevin damping times $\tau_{\rm T}$. At small damping times– where stiff vibrational modes are critically damped– they equilibrate rapidly. However, this also leads to a slower overall relaxation of $C_{11}$. In contrast, larger damping times result in slower equilibration of high–frequency modes, but the average value around which they fluctuate converges more quickly to the long–time relaxed value of $C_{11}$. A trend that is also evident for the simplified model in Figure 2.
@@ IV.3 Applications
For the production runs of all systems studied in this work, we selected damping times, $\tau_{\rm T}$, that ensure rapid convergence of the elastic constants, $C_{ij}$, to their reference values. Added advantage of the single damping time scheme is that it is easy to implement even in the complex polymeric systems. Figure 8 summarizes the time evolution and convergence behavior of $C_{ij}$ for each system. Notably, after the initial rapid relaxation, the amorphous samples exhibit a secondary, slower decay in the elastic constants $C_{ij}$ over time. This effect becomes increasingly pronounced with greater structural complexity– for instance, the decay is noticeably slower in amorphous cellulose compared to amorphous silicon (see Figures 8e, k, and n). A similar two–stage relaxation trend is observed in the case of the generic diatomic harmonic chain with thermal fluctuations (see Figure 1). At first glance, and in the absence of the theoretical insights provided by Figure 1, this transient behavior might be misinterpreted as an indication of flow or irreversible deformation within the sample. To rule out this possibility, we conducted a series of additional simulations involving energy minimization on pre–strained cellulose samples. In these tests, the atomic coordinates remained unchanged after relaxation to the next energy minimum between the initial and final configurations (data not shown), thereby confirming the absence of any flow–like behavior. Taken together, these observations suggest that the slow decay in $C_{ij}$ is not due to structural rearrangements but rather arises from intrinsic vibrational eigenfrequencies of the system. Specifically, the broad distribution of eigenfrequencies characteristic of polymeric and amorphous materials leads to a slow convergence of the elastic response. In particular, the overpopulation of low–frequency vibrational modes– commonly referred to as the Boson peak [41, 42]– is a key factor underlying the long–time decay of $C_{ij}$.
Table 1: Average values of the elastic modulus components. They are computed by block–averaging over ten blocks taken from the second half of the data shown in Figure 8. The standard errors are calculated across these ten blocks. For comparison, available literature data for different systems are also included. In all quoted experimental values, $G$ was obtained from $G=\left(C_{11}-C_{12}\right)/2$.
@@
Argon This work $0.868\pm 0.001$ $4.069\pm 0.002$ $2.325\pm 0.001$
@@
Experiment [43] 1.0 4.0 2.0
Silicon This work $42.80\pm 0.03$ $154.93\pm 0.04$ $69.42\pm 0.02$
@@
Silicon This work $28.1\pm 0.2$ $124.1\pm 0.1$ $66.5\pm 0.1$
@@
Experiment [47] 48.8 156 58.4
PMMA This work $1.85\pm 0.07$ $6.37\pm 0.06$ $3.15\pm 0.07$
@@
Cellulose This work $2.71\pm 0.03$ $14.84\pm 0.02$ $8.62\pm 0.03$
Cellulose Acetate This work $2.09\pm 0.01$ $10.85\pm 0.02$ $6.59\pm 0.01$
A summary of the averaged $C_{ij}$ values, along with available literature data for comparison, is provided in Supplementary Table 1. Overall, the results exhibit good agreement with previously published values, thereby supporting the consistency and reliability of our computational approach. In the case of silicon, our estimates align much better with previous simulations [46] than with experimental results [47], where discrepancies are particularly large for the shear modulus of the amorphous solid. We believe that this is mostly caused by deficiencies of the potential rather than by our comparatively large quench rates, which produce less stable structures than slower cooling protocols. We come to this conclusion, because changing the quench rate by an order of magnitude barely changes $G$, as shown in the Supplementary Section S1 and Supplementary Figs. S1 and S2. Moreover, estimates of $C_{11}$ are already off by 10 GPa for the crystal, which has a well-defined reference structure. Thus, force-field induced errors of order 10 GPa in all elastic tensor elements must be expected. However, what matters most for our work, in which we aim to establish a lean method for the computation of elastic constants, is that previous simulations [46] using similar potentials find similar values as we do. Analyzing the shortcoming of the potential for a covalently bonded solid and suggesting how to mitigate deficiencies is beyond the scope of this work. Since non-covalent interaction, which primarily dictate the elastic tensor coefficients in polymeric systems, have a much reduced many-body character compared to interactions in covalently bonded solids, we expect our results on polymers to better compare to real systems than for silicon.
Notably, the values reported for cellulose and cellulose acetate are predictive in nature, as– to the best of our knowledge– no corresponding $C_{ij}$ data are currently available in the simulation or experimental literature. These predictions provide a valuable reference point for future studies. Furthermore, the computed $C_{ij}$ values for cellulose derivatives across a range of temperatures have already been successfully employed to estimate thermal transport coefficients $\kappa$ in related work [36], which was the main motivation behind the development of the noise-cancellation approach proposed herein. Lastly, we emphasize that kinetic-energy contribution to the computed $C_{ij}$ values are negligible in all cases, remaining below 0.1%, and therefore do not influence the reported results.